One of two complementary angles is 20 degrees larger
than the other. Write and solve an equation that
represents this situation. Find the measure of
each angle.

Answer by ikleyn(52797)

.

    x + (x+20) = 90

    2x         = 90 - 20

    2x         =   70

    x          =   70/2 = 35.


ANSWER. "One" angle is 35 degrees. Another angle is  35 + 20 = 55 degrees.

Solved.